OpenStudy (anonymous):

A sum of money at simple interest amounts to `5852 in 3 years and `7788 in 7 years. What is the rate of interest per annum?

OpenStudy (anonymous):

how to find principal

OpenStudy (anonymous):

formula okay.simpleinterest=p*r*t/100

OpenStudy (perl):

yes thats correct

OpenStudy (anonymous):

principal amount

OpenStudy (perl):

actually we need a different formula

OpenStudy (perl):

Amount = Principal + Interest Amount = P + P*r*t

OpenStudy (perl):

thats why it wasn't working :)

OpenStudy (anonymous):

actually i posted a wrong question sorry

OpenStudy (perl):

lets simplify that formula a bit A = P + P*r*t A = P ( 1 + r*t) A sum of money at simple interest amounts to `5852 in 3 years and `7788 in 7 years. What is the rate of interest per annum? 5852 = P ( 1 + r*3) 7788 = P ( 1 + r * 7 )

OpenStudy (perl):

ok

OpenStudy (anonymous):

yes

OpenStudy (perl):

did you want to solve that one ?

OpenStudy (perl):

it is a system of equations in two unknowns

OpenStudy (anonymous):

yes .i am searching for that question.i forgot

OpenStudy (anonymous):

An amount of money grows up to Rs 11520 in 2 years and up to Rs 13824 in 3 years on compound interest. What is the sum?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

is it continuous compound interest?

OpenStudy (anonymous):

c.I=p(1+r/100)^n

OpenStudy (anonymous):

i dont know what you mean

OpenStudy (perl):

nevermind, there is something called continuous compound interest

OpenStudy (anonymous):

okay.

OpenStudy (perl):

An amount of money grows up to Rs 11520 in 2 years and up to Rs 13824 in 3 years on compound interest. What is the sum? Using your formula, substituting, we have 11520 = P ( 1+ r/100 ) ^2 13824 = P ( 1 + r / 100) ^3

OpenStudy (perl):

notice that they both equal to P 11520/ ( 1 + r / 100) ^2 = P 13824/ ( 1 + r / 100) ^3 = P

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

11520=13824/((100+r)/100)

OpenStudy (anonymous):

we have to find r ah?

OpenStudy (perl):

yes we can solve for r

OpenStudy (anonymous):

1200/100+r

OpenStudy (perl):

11520/ ( 1 + r / 100) ^2 = P 13824/ ( 1 + r / 100) ^3 = P since they both equal to P they equal to each other 11520/ ( 1 + r / 100) ^2 = 13824/ ( 1 + r / 100) ^3 so we have to solve the last equation

OpenStudy (perl):

ok so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

11520=13824/((100+r)/100)

OpenStudy (perl):

ok looks like you did that part correct

OpenStudy (perl):

(sorry, catching up )

OpenStudy (perl):